NAOJ GW Elog Logbook 3.2

I have calibrated the PSD used for the optical lever of the filter cavity mirrors. PSD model : PDP90A thorlab (see file attached).
I computed the calibration for 4 different values of the power (V_SUM should not exceed 4 V)
V_SUM | Normalised_calibration |
1.75 V | 0.00554 m |
2.08 V | 0.00542 m |
3.3 V | 0.00539 m |
3.96 V | 0.00536 m |
Mean = 0.00543 m std = 0.00008 m
Normalised calibration seems reasonably independent of the power.
In order to recover the appropiate calibration ( m/V) to convert a voltage signal into a displacement of the beam on the PSD, this value shoud be dived by the measured V_SUM.

After closing the control loops of pitch and yaw of the telescope mirror in the PR tank, I tried to calibrated the signal in order to have an estimation of the angular motion of the mirror.
The displacement of the beam on the PSD given to a rotation in yaw on of an angle delta is given by
x = 2* arm* delta
Where arm is the distance between the mirror and the PSD.
A dispacement of the beam on the PSD of an amount x corresponds to a PSD voltage output of x/cal.
So delta = Vout*cal /(2*arm)
I measured the calibration factor of the PSD for 4 different powers in order to check if the normalized calibration (that is the calibration divided by the PSD voltage sum) was constant.
I found
V_sum | Normalized calibration |
6.94 V | 0.0071 m |
11.5 V | 0.0068 m |
13.6 V | 0.0073 m |
14.4 V | 0.0071 m |
mean (n_cal) = 0.0071 std (n_cal) = 0.0002 In order to recover the appropriate calibration (m/V) this value should be dived by the V_sum measured each time.
Assuming arm = 1.20 m +/- 0.15 m (the precison on this measurement can be increased by measuring the optical lever arm next time we open PR tank)
and having measured V_sum = 13.3 V
we have
Cal_tot = n_cal/(2*arm*V_sum) = (2.2 +/- 0.3 ) e-4 [1/V] (percentage error 13% to be improved by better measuring the arm)
The comparison between the calibrated spectra with open and closed loops for pitch and yaw is shown in figures 1 and 2.
Some remarks:
1)In both cases the RMS seems to be dominate by the displacement in the region between 3-10 Hz.
2) Since the optical lever makes use of two steering mirrors directly fixed on the stack (see entry 262), this could not be a real motion of the mirror but a motion of the stack
3) This shoud be understood in order to improove the filter shape (Shoud we gain in that region or not?)
NB. For the pitch calibration we need to take into account an additional factor, equal to the the cosine of the incidence angle of the beam on the mirror (see 3rd attache picture)
y = 2* arm* delta* cos(alpha)
THIS FACTOR HAS NOT BEEN TAKEN INTO ACCOUNT IN THE PITCH PLOT WHICH SHOULD HAVE BEEN MULTIPLIED BY A FACTOR 1.412 (since alpha is 45°)

I ckecked the dimensions of BS intermediate mass.
I found that clamp parts are common for NM, EM ans BS itermediate masses.
Therefore, I can conclude that a hanging jig for NM and EM mirror can use for BS.

Since the fans of the booth have been off for a couple of months. I had to clean everything from the dust. I started from the top shelves, wiping one by one all the objects. I moved the boxes made of paperboard out of the clean booth because paperboard is known to produce dust. I cleaned the optical table and all the objects on it. Wiping with a wet tissue was not enough because the tissue releases fibers and dust. So I used the strong green lamp to watch the dust particles, the spray air to blow on the surfaces to make the dust fly and the vacuum cleaner to blow it up from the air. The vacuum cleaner was outside, I only brought the pipe inside. After that, I measured again the particles number.
Size |
cumulative |
counts/L |
0.3um | 664 | 23 |
0.5um | 557 | 19 |
1um | 452 | 16 |
2um | 384 | 12 |
5um | 260 | 9 |
10um | 194 | 7 |
AC | DC | AC/DC | |
no sample | 14uV | 5.9V | 2.5e-6 |
small sapphire sample | 11uV | 5.2V | 2e-6 |
where the AC value is calculated as sqrt(std(X)^2+std(Y)^2)

This afternoon I measured the spectra and the transfer functions of the mirror installed on in the PR tank (to be used as a part of the injection telescope of the filter cavity).
The mirror motion in pich and yaw is sensed by means of a optical lever. The spectrum of the motion in the two degree of freedom is shown in fig1 and 2 of the attached file.
In order to measure the transfer functions, I injected white noise in pitch and yaw (with an amplitude of 3 V ). To improve the diagonalization of the sensing, I changed the the sensing matrix appying a rotation of 0.04 rad. I have also slightly changed the driving matrix, to reduce the excitation of the yaw resonance when injecting noise on pitch. The comparison between the TFs before and after these changes are shown in fig 3-4 and 5-6 for yaw and pitch respectively.

When we install the BS suspension,
(1) one magnet came off, and
(2) the lower suspsnsion wires were broken.

Since one possible explanation of the noise is dust, I used the particle counter to quantify the dust.
Instrument name: MET ONE Airborne Particle Counter HHPC6+
Acquisition time: 10 minutes
Volume: 28.36L
Inside the clean booth of Tama central room.
Size |
cumulative |
counts/L |
0.3um | 199489 | 7034 |
0.5um | 15086 | 532 |
1um | 2018 | 71 |
2um | 740 | 26 |
5um | 114 | 4 |
10um | 57 | 2 |
Almost the same order of a normal room. Indeed I realized that the fans of the clean booth were OFF since a couple of month ago, when I checked whether the acoustic noise was important.
I switched ON the fan and wait half an hour
Size |
cumulative |
counts/L |
0.3um | 943 | 33 |
0.5um | 65 | 2.3 |
1um4 | 4 | 0.14 |
2um | 1 | 0.03 |
5um | 0 | 0 |
10um | 0 | 0 |
After one hour
Size |
cumulative |
counts/L |
0.3um | 139 | 5 |
0.5um | 16 | 0.5 |
1um | 1 | 0.03 |
2um | 1 | 0.03 |
5um | 0 | 0 |
10um | 0 | 0 |
The morning after I came inside the clean booth and run the count
Size |
cumulative |
counts/L |
0.3um | 1018 | 36 |
0.5um | 83 | 3 |
1um | 3 | 0.1 |
2um | 2 | 0.07 |
5um | 0 | 0 |
10um | 0 | 0 |
So it looks that an entire night doesn't clean the air much more than an hour.
Moving very gently I repeated the count
Size |
cumulative |
counts/L |
0.3um | 195 | 7 |
0.5um | 21 | 1 |
1um | 2 | 0.07 |
2um | 0 | 0 |
5um | 0 | 0 |
10um | 0 | 0 |
I think there is some dust on the everything, so when I move the air, the dust flies and make the count higher.
Then I moved the particle counter in the small clean booth on the small optical table that is used for gluing work
Size |
cumulative |
counts/L |
0.3um | 2 | 0.07 |
0.5um | 0 | 0 |
1um | 0 | 0 |
2um | 0 | 0 |
5um | 0 | 0 |
10um | 0 | 0 |
Very clean...
And then I went to clean room of ATC
Size |
cumulative |
counts/L |
0.3um | 6 | 0.2 |
0.5um | 0 | 0 |
1um | 0 | 0 |
2um | 0 | 0 |
5um | 0 | 0 |
10um | 0 | 0 |

I got some information about TAMA Faraday Isolator from Takahashi-san.
I upload these files here.

In the past days I have implemented the control of the end mirror of the filter cavity, following what was already done for the input mirror.
The transfer functions of the mirror motion, measured injecting white noise (with amplitude 3 V) in each degree of freedom, are plotted in the first figures of the attached document. In figure 4,open loop transfer functions are shown. In the last figures the comparison between the spectra with closed and open loops has been plotted.

I made a scan of the bulk reference sample for many positions of the detection unit as I did in this post for other samples.
The position closest to the sample is 34mm, which is the usual position. Other positions are gradually further from the sample.
Modulation reference is from the chopper at 430Hz. Pump power is 30mW before the sample.
Plot1 shows the scan of the sample for different positions of the detection unit. AC signal
Plot2 shows the scan of the sample for different positions of the detection unit. AC signal / DC
For each scan, I took the point at which the calibration value is taken (3rd mm of the scan) and I took 10 minutes of data to check how the noise looks like with a large signal.
Plot3 shows the calibration signal and noise (Y/DC vs X/DC) for each detection unit position.
Plot4 is a zoom of Plot3
The following table is a summary of the values.
position | AC | DC | Phase | R |
(mm) | (V) | (V) | (°) | (W-1) |
34 | 0.12368±0.00087 | 4.87±0.02 | 115.70±0.09 | 0.728±0.006 |
30 | 0.1124±0.0015 | 4.24±0.02 | 113.35±0.11 | 0.76±0.01 |
25 | 0.095±0.001 | 3.78±0.02 | 112.21±0.12 | 0.725±0.009 |
20 | 0.089±0.001 | 3.31±0.02 | 109.76±0.12 | 0.77±0.01 |
15 | 0.0983±0.0008 | 3.56±0.02 | 109.1±0.1 | 0.792±0.008 |
10 | 0.0538±0.0006 | 2.05±0.02 | 109.01±0.12 | 0.75±0.01 |
Error values are the size of the clouds of points on the XY plane (standard deviations)
I notice that
- the calibration factor (R=AC/DC/abs/Power) doesn't change more than 10% when moving the detection unit and doesn't show a clear trend
- the noise on the XY plane is more on the AC value rather than on the Phase value (in other words the cloud is squeezed )
- both the AC and DC get smaller when putting the detection unit further.
- the measurement at 10mm is not reliable because the reflection of the probe on the prism after the sample was on the boundary of the prism.

I placed back the magnetic translation stage and measure again the noise of the small sapphire sample.
I took the measurement in two cases: with and without the small sapphire sample.
Acquisition time: 10 minutes, Rate 100ms, demodulation with the lockin internal oscillator frequency 420Hz.
Plot1 shows the 2D plot of the AC signal from the lockin after demodulation, divided by the DC.
Plot2 is a zooming of the first plot, the circles are centered on the mean of the signal over all the acquisition time and have a radius equal to the standard deviation.
It looks the system is noisier without the sample, this makes me think that vibrations of the sample don't make a lot of difference. I have the idea that most of the noise comes from the dust, and it depends on how we move the air during replacing of the sample or working inside the box.

The signal that I get with the oscilloscope comes from Lock-in CH1 OUTPUT. It is much higher than the signal recorded by the vi (the AC signal in entry 252, for example).
I read the sr830 lock-in manual and I found that the CH1 OUTPUT voltage is proportional to the AC signal according to the following formula:
Output = (signal/sensitivity - offset) x Expand x 10 V
The Expand factor is 1, the sensitivity is 1mV, as we can see in the picture of the front panel.
So in the case of signal = 15uV , for example, I get Output = 150mV, a factor 10^4 higher

[Manuel, Tatsumi]
We checked the wire connections for the coils. The pictures show the order of the wire connections on the suspension, and outside the tank.

[WORKERS] Tatsumi, Takahashi, Manuel, Eleonora
(1) At TAMA south end room (EM2 tank)
* Install four coils.
* Connect in-vacuum cables for the coil
Manual checked the cable connections. He will report soon.
(2) At TAMA center room (BS tank)
* Open the BS tank
* Remove the suspension with BS mirror
* Close the tank
Tatsumi will glue magnets on BS mirror in the next week.
And then we will install the mirror to the BS tank.

I calculated the beam size of the probe beam using OSCAR.
I used the distances I measured and summarized in the first drawing.
The plots show the beam waist along the optical path, cyan area is the sample, vertical lines are the optical component of the experiment, black line is a mirror, cyan line is the f=50mm lens, orange line is the small sphere f=1.25mm
The last image is a comparison of the beam spot size at the PD position with the PD size in the 3 cases.

[Eleonora, Manuel, Tatsumi, Raffaele]
We took two viewport shelves from NM1 tank and installed them at the North-West and South-West viewports of PR tank.
We placed a laser and a mirror on the North-West viewport shelf and placed a PSD on the South-West viewport shelf.
We tried to send the laser to the front surface of the mirror. Since the mirror suspension is not centered on the stack, there is not enough space to get the reflected beam on the other viewport. We decided to send the laser on the back surface of the mirror.
We placed to mirrors as show in the picture. Part of the laser is transmitted and goes to hit the tank wall. Part of the transmitted beam is reflected by the second surface and goes to hit the tank wall on the other side.
We closed the tank.
Then we checked the T, X and Y signals of the PSD using an oscilloscope. The optical lever looks working fine. We remark that the X signal (Yaw motion) is strongly dominated by an oscillation at about 1Hz.

I installed a viewport on the flanges of the cryostat to test its vacuum compatibility. Right now, it looks quite good. The turbo pumb is working well and the viewport seems to be fine. I will leave the system on over the weekend to see whether we can reach the target pressure of 4*10^(-4) Pa.

(1) Wiring parts
See attached picture.
Product Name | Manufacturer | Product No. |
Burndy Male pin | COSMOTEC | PAC16 |
Socket contact (Large, Female) | COSMOTEC | SVC24 |
BNC socket | Custom made | --- |
BNC connector | KYOCERA | BNC-R-F |
(2) Coil support plates
Drawing No. | Length | Qty. | Will use for EM2 | |
Side | F14 | 53 mm | 3 | 2 |
Upper | F15-1 | 43 mm | 6 | 1 |
Lower | F15-2 | 38 mm | 2 | 1 |

Given the fact that a thick sample changes the optical path of the probe, I wanted to see how and why the noise level changes when I change the position of the detection unit. The detection unit is made by one flat mirror at 45°, a f=50mm lens, a reflecting sphere f=2.5mm, and the photodetector.
I turned OFF the chopper to avoid any possible vibration, I set the lock-in internal oscillator as reference frequency (demodulation) at 420Hz.
I connected the oscilloscope at the photodetector (to see the DC signal), and at the output of the lock-in amp (to see the AC signal *1e6). I took some quick measurements, for different positions of the detection unit. The DC has a repeatability of 0.2V, the AC measurement is very rough, just an average of the signal in 10s. Every time I moved the detection unit I had to realign the beam on the PD, tuning the position of the 50mm lens to maximize the DC. The position can be changed only by 35mm, the length of the micrometer screw.
In the following table, there are the DC and AC values at different positions and for different samples. A higher position value means the detection unit is closer to the sample. Hence, 0mm is the furthest point where I could place the detection unit. To move it more it's necessary to unscrew the unit from the board.
I used three samples: the Sapphire small sample diam 1.5" x 5mm, the Sapphire Tama-sized sample diam100mm x 60mm, and the glass KAGRA-sized sample.
No sample | Small sample | Tama-size sample | KAGRA size sample | |||||||||
Position | DC | AC | DC | AC | DC | AC | DC | AC | ||||
(mm) | (V) | (mV) | (V) | (mV) | (V) | (mV) | (V) | (mV) | ||||
34 | 7.4 | ~100 | 6.9 | ~140 | 8.5 | ~500 | 8.6 | ~1100 | ||||
30 | 6.8 | 6.4 | ~100 | 8.5 | ~500 | 8.6 | ~900 | |||||
25 | 5.8 | ~80 | 5.4 | ~110 | 8.5 | ~400 | 8.6 | ~800 | ||||
20 | 4.9 | 4.6 | ~90 | 8.0 | ~150 | 8.6 | ~900 | |||||
15 | 4.4 | ~80 | 4.0 | ~100 | 7.4 | ~120 | 8.6 | ~900 | ||||
10 | 4.0 | 3.6 | ~90 | 6.6 | ~115 | 8.5 | ~600 | |||||
5 | 4.4* | 3.6 | ~80 | 5.9 | ~80 | 8.5 | ~400 |
Looking at those data, I can say:
- the DC decreases when the unit is placed further. This is reasonable considering the finite size of the PD and the divergence of the beam.
- In the case without any sample, when the unit position changes, the AC noise level doesn't change a lot.
- In the case with the small sample, when the unit distance increases, the AC noise level does decrease, maybe proportionally to the DC.
- In the case with the Tama-size sample, when the unit distance increases, the AC noise level changes a lot, compared to the DC.
- In the case with the KAGRA-size sample, when the unit distance increases, the DC is pretty constant and the AC noise level changes a bit.
The hypothesis I have in mind is that the probe spot size makes an important role when compared to the detector size.
I will use my simulations to try to reproduce the behavior shown in those measurements and try to find an explanation.
The signal that I get with the oscilloscope comes from Lock-in CH1 OUTPUT. It is much higher than the signal recorded by the vi (the AC signal in entry 252, for example).
I read the sr830 lock-in manual and I found that the CH1 OUTPUT voltage is proportional to the AC signal according to the following formula:
Output = (signal/sensitivity - offset) x Expand x 10 V
The Expand factor is 1, the sensitivity is 1mV, as we can see in the picture of the front panel.
So in the case of signal = 15uV , for example, I get Output = 150mV, a factor 10^4 higher

Aoyama-san of National Institute of Polar Research warmed up the Iodine stabilized He-Ne laser.
But they found trouble on the laser.
Now we are waiting for the warming-up run
and for the stable operation.
We will check the laser in tomorrow morning.